OpenStudy (anonymous):

how to write an equation in slope intercept form for a line that passes through the given point and is parallel to the given line y=5

OpenStudy (campbell_st):

well the line y = 5 has a slope of zero... the general slope intercept form is y = mx + b m = slope and b = y intercept... so if m = 0 substitute it to find the equation of the line

OpenStudy (anonymous):

given parallel to y , so anyline parallel to the the line y=5 and slope as m has the form y=mx+c here m=0 since, y= mx+c since m=0 then y= (that point y-intercept) enjoy ;)